OpenStudy (anonymous):

y=√4x-13 find the domain can somebody explain step by step please

jimthompson5910 (jim_thompson5910):

how far does the square root extend? is it \(\Large y = \sqrt{4x-13}\) ?? OR is it \(\Large y = \sqrt{4x}-13\) ??

OpenStudy (anonymous):

the first one

jimthompson5910 (jim_thompson5910):

You cannot have the terms under the square root be negative, so we have to force 4x-13 to be 0 or positive which means you need to solve this for x \[\Large 4x - 13 \ge 0\]

jimthompson5910 (jim_thompson5910):

what do you get when you do so?

OpenStudy (anonymous):

\[4x \ge 13\]

jimthompson5910 (jim_thompson5910):

you can then divide both sides by 4 to get \[\Large x \ge \frac{13}{4}\]

jimthompson5910 (jim_thompson5910):

so the domain is the set of all real numbers x such that \(\Large x \ge \frac{13}{4}\)